# CoDial: Interpretable Task-Oriented Dialogue Systems Through Dialogue Flow Alignment

### v1

@@ 6 Conclusion
In this work, we introduced CoDial, a novel framework for building interpretable TOD systems by grounding structured dialogue flows to programmatic guardrails. CoDial introduces CHIEF, a heterogeneous graph representation of dialogue flows, and employs LLM-based code generation to automatically convert dialogue flows into executable guardrail specifications (e.g., in NVIDIA’s Colang), enabling zero-shot creation of interpretable TOD systems with minimal expert effort. Moreover, through iterative manual and LLM-aided refinements, CoDial supports rapid incorporation of domain-expert feedback, further enhancing the generated code. Our empirical findings support CoDial’s effectiveness, achieving SOTA performance among inference-based methods on STAR and competitive results on MultiWOZ.
@@
While CoDial offers an interpretable and modifiable approach to TOD systems, it has certain limitations. First, scalability remains a challenge. For large and complex dialogue flows, CoDial re-queries all slots every turn, which may increase latency and computational cost. In general, the DST is the biggest challenge of the system and we leave improvements to future work. Second, CoDial is less effective for multi-domain dialogues, as it operates on a single dialogue flow at a time. Handling seamless transitions between multiple domains would require additional mechanisms beyond the current framework, and structured logic on how different domains flow into each other, which we leave to future work. In general, dialogue flows assume there is a logical progression to conversations, which isn’t the case for all dialogue structures. Finally, reproducibility could be a concern when relied on API-based LLMs. Since these models are periodically updated, responses may vary across different API versions, making consistent evaluation challenging.
@@ Appendix A Appendix
A.1 Details on Code Generation
@@
We represent the graphs defined by CHIEF (Section 3.1) as text in JSON format. The JSON representation consists of a list of nodes and a list of edges. The node list defines the dialogue flow nodes, specifying their types and assigning each a unique identifier (node ID). The edge list specifies the connections between nodes using their IDs (Figure 8). The JSON nodes, global and fallback actions, and functional specifications for function calls are translated into Colang code with our automatic code generation pipeline. The external action node functions, referred to as Actions in Colang, are implemented in Python.
@@
Figure 4: Execution life cycle of the generated agent.
@@
The algorithm designed in Colang (Algorithm 1) determines whether a request node should be executed (i.e., prompt the user for information) by checking the values of its associated slots. To track the state of other node types, we instruct $\text{LLM}_{\text{GCG}}$ to define helper variables following a structured naming pattern, where <id> represents the corresponding node’s ID:
@@
Figures 7 and 8 present an example of the STAR task schema pictures, our JSON representation of the corresponding dialogue flow, and the generated Colang code.
A.2 STAR Implementation Details
@@
We convert the STAR task schemas, originally provided as images, into CHIEF representation described in Section 3.1. We use one-shot prompting with GPT-4o to convert pictures into JSON. We convert yellow nodes in pictures into conditions for edges. However, we observed that GPT-4o occasionally misassigns edge connections, requiring manual corrections. Additionally, we enrich the JSON representations by adding more context, such as example values for each slot. We also define hello action as the only global action and goodbye, out_of_scope, and anything_else as fallback actions for all tasks.
@@
For evaluation, since we are working with offline conversations (i.e., the user is not interacting with the actual TOD system), we approximate the wizard’s state at the end of each turn and adjust the program’s state accordingly. This helps prevent the program’s state from deviating from the ground-truth conversation. To achieve this, we first find the node in dialogue flow that the ground-truth conversation was in by mapping the ground-truth action label, if available, to a node in the dialogue flow. We manually create this mapping from action labels to the dialogue flow nodes. Next, we use depth-first search to trace the path from the start of the dialogue flow to the current conversation node. Finally, we adjust each state variable based on whether the corresponding node is part of the current conversation pathway, as described in Algorithm 2.
@@
3:$n_{tgt}\leftarrow\phi(a_{gt})$
4:$n_{v}\leftarrow v.node$
5:$P\leftarrow$ DFSPath($G,G.start,n_{tgt}$)
@@
A.3 MultiWOZ Implementation Details
@@
1:Training set $\mathcal{D}_{\text{train}}$, Validation set $\mathcal{D}_{\text{val}}$, Instruction $I$, Agent $\text{LLM}_{\text{A}}$, Optimizer LLM $M$, Batch size $B$
2:$\hat{Y}_{\text{val}}\leftarrow$ DST($\mathcal{D}_{\text{val}}.H,\text{LLM}_{\text{A}},I$)
@@
Our optimization algorithm is summarized in Algorithm 3. For each DST variable $v^{(s)}_{j}$, we randomly sample two mutually exclusive sets of conversation turns to serve as training and validation sets. The training examples are divided into batches of 5, and each batch is used to guide the optimizer GPT-4o model to rewrite the instruction $p^{(s)}_{j}$, resulting in a candidate prompt. If the revised instruction improves performance on the validation set, it is retained; otherwise, the original is kept, ensuring that modifications are only accepted when they lead to measurable improvements.
@@
A.4 Experimental Details
For the STAR dataset, we compute BLEU-4 score (Papineni et al., 2002). using SacreBLEU (Post, 2018). We also follow Mosig et al. (2020) to compute F1 and accuracy. Since we applied STAR’s response templates for response generation, we use regex patterns to match generated responses with actual values to a template. For the MultiWOZ dataset, we compute BLEU, Inform and Success rates, and Joint Goal Accuracy (JGA) using the official evaluation script (Nekvinda and Dušek, 2021). We report the mean result of three runs.
If a generated program contains syntax or runtime errors, we regenerate the code to obtain a functional version. The only exception is Gemini 2.0 Flash, which struggles with calling our defined Colang helper flows. Since this issue is minor, we manually correct the syntax to assess the model’s ability to generate programmatic logic for dialogue flows. We access OpenAI models through OpenAI and other models through OpenRouter [^10] https://openrouter.ai/ API.
@@ A.4 Experimental Details
We modify NeMo’s default value_from_instruction prompt structure to begin with a system message, followed by the entire conversation history and instructions combined into a single user message (Figure 5). During our initial experiments, we suspected that NeMo’s original prompt structure—where each message in the conversation history was passed as a separate user or assistant message—hindered $\text{LLM}_{\text{A}}$’s ability to follow instructions effectively.
@@
Figure 7: Example of a generated code for STAR Ride Change task
@@ A.5 Detailed Baselines
  • SGP-TOD (Zhang et al., 2023) is a purely generative approach that uses two-stage prompting to track dialogue state and generate response. It employs graph-based dialogue flows to steer LLM actions, ensuring adherence to predefined task policies without requiring fine-tuning or training data.
    To ensure a fair comparison, we replicated their setup using the same newer $\text{LLM}_{\text{A}}$ model as ours (Table 7). We ran their released code without modification, except for switching the API model to GPT-4o-mini. Surprisingly, performance dropped significantly. After contacting the authors, they advised adapting the prompt structure to the aligned LLMs—placing instructions in the system message and including examples and dialogue history in the user message. However, even with this adaptation, the performance did not match the results originally reported with GPT-3.5, suggesting that a generative approach could not be a trivial solution and requires careful prompt engineering. Figure 6 further illustrates differences between CoDial and SGP-TOD through two cherry-picked examples.

### v2

@@ 6 Conclusion
In this work, we introduced CoDial, a novel framework for building interpretable TOD systems by grounding structured dialogue flows to programmatic guardrails. CoDial introduces CHIEF, a heterogeneous graph representation of dialogue flows, and employs LLM-based code generation to automatically convert dialogue flows into executable guardrail specifications (e.g., in NVIDIA’s Colang), enabling zero-shot creation of interpretable TOD systems with minimal human effort. Moreover, through manual and LLM‐aided refinements, CoDial supports rapid incorporation of user feedback, further enhancing the generated code. Our empirical findings support CoDial’s effectiveness, achieving SOTA performance on STAR and competitive results on MultiWOZ in a strict zero-shot setting.
@@
While CoDial offers an interpretable and modifiable approach to TOD systems, it has certain limitations. First, scalability remains a challenge. For large and complex dialogue flows, CoDial re-queries all slots every turn, which may increase latency and computational cost. In general, improving DST performance and efficiency remains a potential direction for future work. Second, CoDial is less effective for multi-domain dialogues, as it operates on a single dialogue flow at a time. Handling seamless transitions between multiple domains would require additional mechanisms beyond the current framework
Moreover, developing measurable metrics for interpretability and user accessibility, two central motivations of this work, remains an open direction. We demonstrate interpretability indirectly through modifiability, as shown in the incorporation of human feedback in CHF. In terms of accessibility, an ideal study would evaluate the effort required for users to represent their knowledge as a task schema (e.g., CHIEF representation in CoDial) and compare it across approaches. While CoDial abstracts away manual programming, certain applications may still require some familiarity with LLM guardrails and Colang for effective modification. CoDial mitigates this through textual refinement interfaces (RIs), though their adequacy ultimately depends on the specific use case.
@@ Appendix A Appendix
A.1 Detailed Related Work
Task-Oriented Dialogue
Building generalizable conversational systems is challenging due to the complexity of human conversations, particularly when domain expertise is involved (Chen et al., 2017), leading to a focus on task-oriented systems for specific domains (Jacqmin et al., 2022). While LLMs have demonstrated impressive capability in a wide variety of domains, they struggled with TOD and fell behind if not used properly Hudeček and Dusek (2023). Some research (Zhang et al., 2023; Mehri and Eskenazi, 2021) has used a neural schema-guided approach to generalize TOD systems to unseen tasks without interpretability. AnyTOD (Zhao et al., 2023) provided an interpretable neuro-symbolic approach by viewing task schema as a manually-written policy program that controls the dialogue flow. However, beyond the manual coding requirement, AnyTOD also relied on extensive training with highly similar task schemas. As a result, it suffered substantial performance drops when transferred to even slightly different task structures, revealing limited generalizability to unseen tasks.
Guardrails
CoDial employs guardrails to steer LLMs behaviour. Guardrailing aims to enforce human-imposed constraints to control LLMs in the inference time (Dong et al., 2024; Rebedea et al., 2023; Guardrails AI, ). While originating from AI safety, we argue that they can generally be used to define desired behaviour to constrain. Although traditional dialogue management systems, like Google Dialogflow [^4] https://dialogflow.cloud.google.com, allow rigid modelling of dialogue states, they often lack the flexibility to define complex task logic, and it is difficult for a user to further enhance the system. NVIDIA NeMo-Guardrails (Rebedea et al., 2023) is a toolkit that adds programmable guardrails to LLM-based conversational applications without fine-tuning. NeMo-Guardrails employs Colang NVIDIA (2024), a programming language, to establish highly flexible conversational flows and guide LLMs within them. Dong et al. (2024) suggested using neuro-symbolic approaches to guardrail LLMs, where a neural agent (e.g., an LLM) can deal with frequently seen cases, and a symbolic agent can embed human-like cognition through structured knowledge for the rare cases.
Code Generation and Prompt Optimization
A.2 Details on CHIEF and GCG
(a) $\text{prompt}_{\text{GCG}}$ for $\text{CoDial}_{\text{free}}$
(b) $\text{prompt}_{\text{GCG}}$ for $\text{CoDial}_{\text{structured}}$
@@ A.2 Details on CHIEF and GCG
A.2.1 CHIEF
Below, we discuss the main node types and actions in CHIEF.
Request
External Action
Inform (and Confirm)
Global and Fallback Actions
In addition to nodes, CHIEF supports representing global and fallback actions that are not tied to particular dialogue steps. Global actions can be triggered at any point in the dialogue flow (e.g. responding to a greeting). We also define fallback actions, general responses used when no other action is selected (e.g. Sorry, I can’t help with that).
We represent the graphs defined by CHIEF (Section˜3.1) as text in JSON format. The JSON representation consists of a list of nodes and a list of edges. The node list defines the dialogue flow nodes, specifying their types and assigning each a unique identifier (node ID). The edge list specifies the connections between nodes using their IDs (Figure˜10). The JSON nodes, global and fallback actions, and functional specifications for function calls are translated into Colang code with our automatic code generation pipeline. The external action node functions, referred to as Actions in Colang, are implemented in Python.
A.2.2 GCG
Figure 4: Execution life cycle of the generated agent in $\text{CoDial}_{\text{structured}}$.
Dialogue State Tracking (DST)
Since updating a slot may affect the state (e.g. in a search task, modifying the search criteria requires re-executing the search), CoDial needs to identify the helper variables that need to be invalidated when each slot is updated. We instruct $\text{LLM}_{\text{GCG}}$ to list the helper variables of nodes that are reachable from the updated slot in the graph (i.e., nodes that are direct or indirect children of the slot’s request node). These variables are then reset to null or false, depending on their type, when the slot is updated during execution.
@@ A.2.2 GCG
The $\text{CoDial}_{\text{structured}}$ algorithm designed in Colang (Algorithm˜1) determines whether a request node should be executed (i.e., prompt the user for information) by checking the values of its associated slots. To track the state of other node types, we instruct $\text{LLM}_{\text{GCG}}$ to define helper variables following a structured naming pattern, where <id> represents the corresponding node’s ID:
@@
Figures˜8, 9 and 10 present an example of the STAR task schema pictures, our JSON representation of the corresponding dialogue flow, and the generated Colang code.
A.3 STAR Implementation Details
@@
We convert the STAR task schemas, originally provided as images, into CHIEF representation described in Section˜3.1. We use one-shot prompting with GPT-4o to convert pictures into JSON. We convert yellow nodes in pictures into conditions for edges. However, we observed that GPT-4o occasionally misassigns edge connections, requiring manual corrections. Additionally, we enrich the JSON representations by adding more context, such as example values for each slot. We also define hello action as the only global action and goodbye, out_of_scope, and anything_else as fallback actions for all tasks.
@@
For $\text{CoDial}_{\text{structured}}$ evaluation, since we are working with offline conversations (i.e., the user is not interacting with the actual TOD system), we approximate the wizard’s state at the end of each turn and adjust the program’s state accordingly. This helps prevent the program’s state from deviating from the ground-truth conversation. To achieve this, we first find the node in dialogue flow that the ground-truth conversation was in by mapping the ground-truth action label, if available, to a node in the dialogue flow. We manually create this mapping from action labels to the dialogue flow nodes. Next, we use depth-first search to trace the path from the start of the dialogue flow to the current conversation node. Finally, we adjust each state variable based on whether the corresponding node is part of the current conversation pathway, as described in Algorithm˜2.
@@
[…]
@@
A.4 MultiWOZ Implementation Details
@@
1:Training set $\mathcal{D}_{\text{train}}$, Validation set $\mathcal{D}_{\text{val}}$, Instruction II, Agent $\text{LLM}_{\text{A}}$, Optimizer LLM $M$, Batch size $B$
[…]
@@
Our optimization algorithm is summarized in Algorithm˜3. For each DST variable $v^{(s)}_{j}$, we randomly sample two mutually exclusive sets of conversation turns to serve as training and validation sets. The training examples are divided into batches of 5, and each batch is used to guide the optimizer GPT-4o model to rewrite the instruction $p^{(s)}_{j}$, resulting in a candidate prompt. If the revised instruction improves performance on the validation set, it is retained; otherwise, the original is kept, ensuring that modifications are only accepted when they lead to measurable improvements.
@@
A.5 Experimental Details
We use GPT-4o [^6] https://openai.com/index/hello-gpt-4o/, Claude 3.5 Sonnet [^7] https://www.anthropic.com/news/claude-3-5-sonnet, Gemini 2.0 Flash [^8] https://developers.googleblog.com/en/gemini-2-family-expands/, and DeepSeek V3 (DSV3) (DeepSeek-AI et al., 2024) as $\text{LLM}_{\text{GCG}}$, and GPT-4o-mini, GPT-5, and DSV3 as $\text{LLM}_{\text{A}}$. We access OpenAI models through OpenAI and other models through OpenRouter [^9] https://openrouter.ai/ API  [^10] All models were accessed from January to September 2025..
If a generated program contains syntax or runtime errors, we regenerate the code to obtain a functional version. The only exception is Gemini 2.0 Flash, which struggles with calling our defined Colang helper flows. Since this issue is minor, we manually correct the syntax to assess the model’s ability to generate programmatic logic for dialogue flows.
For the STAR dataset, we compute BLEU-4 score (Papineni et al., 2002). using SacreBLEU (Post, 2018). We also follow Mosig et al. (2020) to compute F1 and accuracy. Since we applied STAR’s response templates for response generation, we use regex patterns to match generated responses with actual values to a template.
For the MultiWOZ dataset, we compute BLEU, Inform and Success rates, and Joint Goal Accuracy (JGA) using the official evaluation script (Nekvinda and Dušek, 2021). Since $\text{CoDial}_{\text{free}}$ does not include an explicit DST component and most MultiWOZ metrics rely on DST predictions, we do not report $\text{CoDial}_{\text{free}}$ results on this dataset.
@@ A.5 Experimental Details
We modify NeMo’s default value_from_instruction prompt structure to begin with a system message, followed by the entire conversation history and instructions combined into a single user message (Figure˜5). During our initial experiments, we suspected that NeMo’s original prompt structure—where each message in the conversation history was passed as a separate user or assistant message—hindered $\text{LLM}_{\text{A}}$’s ability to follow instructions effectively.
@@
Figure 7: Example of code logic in CoDial that enables interpretability. A user can inspect runtime variables to trace the reasoning behind the outputs generated by the TOD system.
Figure 8: Example of a generated code for STAR Ride Change task with $\text{CoDial}_{\text{free}}$
Figure 9: Example of a generated code for STAR Ride Change task with $\text{CoDial}_{\text{structured}}$
@@ A.6 Detailed Baselines
  • SGP-TOD (Zhang et al., 2023) is a purely generative approach that uses two-stage prompting to track dialogue state and generate response. It employs graph-based dialogue flows to steer LLM actions, ensuring adherence to predefined task policies without requiring fine-tuning or training data. To ensure a fair comparison, we replicated their setup using the same newer $\text{LLM}_{\text{A}}$ model as ours (Table˜7). We ran their released code without modification, except for switching the API model to GPT-4o-mini. Surprisingly, performance dropped significantly. After contacting the authors, they advised adapting the prompt structure to the aligned LLMs—placing instructions in the system message and including examples and dialogue history in the user message. However, even with this adaptation, the performance did not match the results originally reported with GPT-3.5, suggesting that a generative approach could not be a trivial solution and requires careful prompt engineering. Figure˜6 further illustrates differences between CoDial and SGP-TOD through two cherry-picked examples. Specifically, in Figure˜6(b), by analyzing the variable values in the runtime, a user can easily spot that the generated output stems from the code snippet in Figure˜7, where it asks for the next missing value, if any.
@@
  • DARD (Gupta et al., 2024) is a multi-agent TOD system that delegates responses across domain-specific agents coordinated by a central dialogue manager. It combines fine-tuned models (Flan-T5-large, Mistral-7B) with large LLMs (Claude Sonnet 3.0), yielding SOTA results on MultiWOZ with significant gains in inform and success rates. However, its performance depends heavily on extensive carefully designed prompt tuning and few-shot examples, limiting efficiency and increasing human effort.